solve each equation if the domain is (-4,-2,0,2,4)
y = -4x-6

Answer by KMST(5328)

This is a very strange question (or so it looks to me). Maybe this is how functions are taught somewhere nowadays.
The domain is the set of x values allowed for a function.
A function assigns one y value to each x value. (The same y value can be assigned to more than one x, but no x can have two different y values in a function).
With only 5 x values in the domain, the function will consist of 5 (x,y) ordered pairs.
For x=-4, y=-4(-4)-6=16-6=10.
For x=-2, y=-4(-2)-6=8-6=2.
For x=0, y=-4(0)-6=0-6=-6.
For x=2, y=-4(2)-6=-8-6=-14.
For x=4, y=-4(4)-6=-16-6=-22.
The function can be represented by the ordered pairs
(-4,10), (-2,2), (0,-6), (2,-14), and (4,-22).
I guess that is what was meant to be the answer.
I do not know what notation was expected.
Maybe the function was expected to be represented as the set
{(-4,10), (-2,2), (0,-6), (2,-14), and (4,-22)}